electric field at a point on the equatorial line of a small electric

dipole is 10N/C. What will be the electric field at a point situated at the

same distance on the axis of the dipole?


1
Expert's answer
2021-06-22T09:47:41-0400

The field at a point on the equatorial line is


Ee=k2pr3.E_e=k\frac{2p}{r^3}.

The field on the axis:


Ea=kpr3.E_a=k\frac {p}{r^3}.

Thus,

Ea=Ee2=5 N/C.E_a=\frac{E_e}{2}=5\text{ N/C}.